Juanita has an apportunity to invest in her friend clothing store.the initial investment is $10,200 and the expected annual cash flow thereafter are ($300,$700,$1,300, $2000,$5000,$5000) . What is Juanita IRR on this investment?
1
Expert's answer
2020-11-17T07:22:58-0500


IRR lies between 9"\%" and 10"\%" . By simple interpolation IRR = "9" "\%" "+" "1" "\%" "\times""\frac{101}{376+101} = 9.21\%"